OpenStudy (anonymous):

a recipe uses 7 cups of flour for 4 loaves of bread. How many cups are needed for 25 loaves of bread?

OpenStudy (anonymous):

7/4 = B/25 Solve for B

OpenStudy (anonymous):

i dont know what B is?

OpenStudy (phi):

This is solved using proportions. In Romero's answer B stands for the number of cups needed for 25 loaves. Do you know how to solve a proportion?

OpenStudy (anonymous):

Nooo i dont /:

OpenStudy (phi):

And here is how to solve the proportion http://www.khanacademy.org/math/algebra/ratios-proportions/v/find-an-unknown-in-a-proportion

OpenStudy (phi):

\[ \frac{7}{4}= \frac{x}{25} \] \[ x= 25\cdot \frac{7}{4} \] \[ x= 43 \frac{3}{4} \]